OpenStudy (anonymous):

WILL AWARD MEDALS NEED FAST HELP WITH A FEW WORD PROBLEMS!!! DESPERATE!! Flying against the jetstream, a jet travels 4200 mi in 4 hours. Flying with the jetstream, the same jet travels 11,790 mi in 9 hours. What is the rate of the jet in still air and what is the rate of the jetstream?

OpenStudy (anonymous):

4200 miles in how many hours?

OpenStudy (anonymous):

You didn't write in the question

OpenStudy (anonymous):

4 hrs

OpenStudy (anonymous):

Fixed it. Sorry

OpenStudy (anonymous):

First, lets start off by taking variables let x be the speed of jet in still air let y be the speed of jetstream also x>y otherwise jet would not be able to fly Now, can you please calculate the speed for both the cases?the distance and time is given \[s=\frac{d}{t}\]

OpenStudy (anonymous):

s = 4200/4 = 1050 s = 11,790/9 = 1310

OpenStudy (anonymous):

ok, good

OpenStudy (anonymous):

x = 1310 & y = 1050. That is really it?

OpenStudy (anonymous):

when flying against the jetstream, the speed of the jetstream offers resistance to the jet, so the speed of jetstream is subtracted from the actual speed of the jet, so speed of jet against stream=speed of jet-speed of jetstream \[x-y=1050\] When flying with the stream, the speeds are added up so, speed with stream=actual speed of jet+speed of jetstream \[x+y=1310\]

OpenStudy (anonymous):

solve these equations to find x and y, can you do that?

OpenStudy (anonymous):

1310 - 1050 = 260 1310+1050 = 2360

OpenStudy (anonymous):

Doesn't feel right

OpenStudy (anonymous):

x is not 1310 and y is not 1050, those are the values we calculated to make an equation

OpenStudy (anonymous):

you are supposed to FIND x and y using those equations

OpenStudy (anonymous):

s = 4200/4 = 1050 s = 11,790/9 = 1310 1310-1050 = 260 1050 + 260 = 1310 1310 - 260 = 1050

OpenStudy (anonymous):

don't put 1310 and 1050 in place of x and y, you don't know those values here's how you're supposed to do it \[x-y+x+y=1050+1310\] notice the lhs of both equations are added, rhs of both equations are added

OpenStudy (anonymous):

2x + 2y = 2360 590

OpenStudy (anonymous):

the y's get cancelled \[2x=1360\]

OpenStudy (anonymous):

2/1680 = 680 x=680

OpenStudy (anonymous):

y = 590 x = 680

OpenStudy (anonymous):

sorry\[2x=2360\]

OpenStudy (anonymous):

2/2360 = 1180 x=1180 y = 590

OpenStudy (anonymous):

\[x+y=1310\]\[y=1310-x\]\[y=1310-1180\]\[y=130\]

OpenStudy (anonymous):

x = 1180 y = 130

OpenStudy (anonymous):

Yep, notice that x>y as I stated earlier

OpenStudy (anonymous):

units are miles/hour

OpenStudy (anonymous):

don't forget to write your units, they are important

OpenStudy (anonymous):

yes. x = 1180 still air y = 130 speed of jet stream

OpenStudy (anonymous):

Thank you 100%!! I have more but don't want to tax you!

OpenStudy (anonymous):